Andy, whose age is n years, is 6 years older than Ricardo. The sum of their age is 38.
Answer by JulietG(1812)

A + R = 38
A = R + 6
Substitute the known value of A in the second equation into the first.
A + R = 38
(R+6) + R = 38
2R + 6 = 38
Subtract 6 from each side
2R = 32
Divide each side by 2
R = 16
.
If Ricardo is 16, then Andy is 22.
22+16 = 38